OpenStudy (anonymous):

Please help!!!!!!!! verify the identity. sin (x+pi/2)=cos x

OpenStudy (anonymous):

please help!!!

Directrix (directrix):

sin (x+pi/2)= sin(x) * cos(pi/2) + cos(x)* sin(pi/2) Crank this out. @z123t6

OpenStudy (anonymous):

I am sorry I dont understand this at all... can you please break it down more

OpenStudy (anonymous):

@directrix

OpenStudy (anonymous):

Use sin (a+b) = sin a . cos b + cos a .sin b

OpenStudy (anonymous):

and thats it?

OpenStudy (anonymous):

Use this formula here: sin (pie/2 + x) ....

OpenStudy (anonymous):

Use \(\sin(\pi/2-x) = \cos(x)\)?

OpenStudy (anonymous):

ahhh i am so confused im sorry i just dont get it

Directrix (directrix):

sin (x+pi/2)= sin(x) * cos(pi/2) + cos(x)* sin(pi/2) sin (x+pi/2)= sin(x) * 1 + cos(x) * 0 sin (x+pi/2)= ? @z123t6

Directrix (directrix):

Trig has a lot of memorizing or use it until you remember it work. In this problem, you have to know the sine of a sum of angles formula. Then, you have to know the sine and cosine values of various angles. Then, you have to use algebra to put all that together.

OpenStudy (anonymous):

oh okay so would sin (x+pi/2) = cos x?

Directrix (directrix):

Based on the work, yes.

OpenStudy (anonymous):

and would that be my final answer or is there more to it?

Directrix (directrix):

This is a proof or demonstration. You have to show all the work so that the answer makes sense. Verify the identity: sin (x+pi/2)=cos x ----------------- sin (x+pi/2)= sin(x) * cos(pi/2) + cos(x)* sin(pi/2) sin (x+pi/2)= sin(x) * 1 + cos(x) * 0 sin (x+pi/2) = cos x

OpenStudy (anonymous):

ohhh okay well thank makes a lot more sense now lol :) thank you so much!

Directrix (directrix):

Glad to help.